# High-dimensional quantum XYZ product codes for biased noise

### v2

@@ III.2 Dimension of 4D XYZ product code
Next, we consider 4D XYZ product of two concatenated codes obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ ($n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd), and compute it dimension. Let the repetition code $C_{1}$ with length $n_{1}$ be the outer code and the code $C_{2}$ with length $n_{2}$ be the inner code, whose parity-check matrices are $H_{1}$ with dimension $\left(n_{1}-1\right)\times n_{1}$ and $H_{2}$ with dimension $\left(n_{2}-1\right)\times n_{2}$ respectively, the corresponding $X$-type and $Z$-type parity-check matrices $H_{x}$ and $H_{z}$ of the concatenated code obtained from $C_{1}$ and $C_{2}$ have the following form:
@@
Corollary 2 (The dimension of 4D XYZ concatenated code).
The dimension $k$ of the code, which is 4D XYZ product of two concatenated codes obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ (where $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd), is 1.
@@
According to Theorem 1 and $\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)=1$, the dimension $k$ of 4D XYZ concatenated code is $1+k_{SV}+k_{TU}$, where $k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ and $k_{TU}=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$.
@@
For a concatenated code $C$ obtained from a pair of repetition codes, whose $X$-type and $Z$-type parity-check matrices are $H_{x}$ and $H_{z}$ respectively, it is easy to prove that $\dim\left(\ker\left(\begin{bmatrix}H_{x}\\ H_{z}\end{bmatrix}\right)\right)=1$ and $\dim\left(\ker\left(\left[H_{z}^{T},H_{x}^{T}\right]\right)\right)=n-1-dim\left(row\left(\begin{bmatrix}H_{x}\\ H_{z}\end{bmatrix}\right)\right)=n-1-(n-1)=0$. The reason is that, for code $C$ the undetectable $Y$-type error only contain one $Y$-type logical operator whose weight is code length. Thus, the dimension $k$ of 4D XYZ concatenated code is 1. ∎
Next, we prove Lemma 1 which will be used in the proof of Corollary 3.
Lemma 1 (The number of independent undetectable $Y$-type errors of 2D toric code).
@@
Since 2D toric code encodes 2 logical operators, there are 2 independent $Y$-type logical operators at most. When the toric code lattice is a square, these two independent $Y$-type logical operators are full diagonal of Pauli $Y$ operators from the upper right to the lower left and from the upper left to the lower right, respectively. In Fig. 4, we give an example of $Y$-type logical operators of 2D toric code with lattice size $3\times 3$.
@@
Any $Y$-type stabilizer must be generated by the product of some $X$-type and $Z$-type stabilizers both of which act on the same set of qubits. According to the topology of toric code, this can only be realized by selecting plaquettes ($Z$-type stabilizers) and vertices ($X$-type stabilizers) along the diagonal of $j\times j$ square lattices periodically. Fig. 5 gives an example of $Y$-type stabilizer of toric code with dimension $j=4,\ k=6$.
Figure 5: A $Y$-type stabilizer of toric code with lattice size $j=4,\ k=6$. The red solid line edges are qubits on which the $Y$-type stabilizer acts. Each plaquette filled with blue-grey is the selected $Z$-type stabilizer, and its top right and bottom left vertices are the selected $X$-type stabilizers.
One can see that when $j$ and $k$ are coprime, namely, $\gcd\left(j,k\right)=1$, we cannot find a selection method of $X$-type and $Z$-type stabilizers to generate a $Y$-type stabilizer, thus the number of $Y$-type stabilizers is zero. While in the case of $\gcd\left(j,k\right)=g\neq 1$, using the selection method as shown in Fig. 5, one can find $2\left(g-1\right)$ independent $Y$-type stabilizers by shifting to left or right, where the factor 2 comes from the left-right symmetry of toric code.
@@
Corollary 3 (The dimension of the 4D Chamon code).
The dimension $k$ of the 4D Chamon code, which is 4D XYZ product of two hypergraph product codes that are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$, is
@@
The hypergraph product of a pair of repetition codes is a toric code[5]. Thus, given two toric codes which are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$, whose parity-check matrixes are $H_{1}$, $H_{2}$, $H_{3}$ and $H_{4}$, respectively, by Definition 1, the corresponding stabilizer matrix s of the 4D XYZ product code is as shown in Eq. (16), where $H_{x_{1}}=\left[I_{n_{1}}\otimes H_{2},H_{1}^{T}\otimes I_{n_{2}}\right]$ with dimension of $m_{1}\times n_{A}$ and $H_{z_{1}}=\left[H_{1}\otimes I_{n_{2}},\ I_{n_{1}}\otimes H_{2}^{T}\right]$ with dimension of $m_{2}\times n_{A}$ are the $X$-type and $Z$-type parity-check matrixes of toric code $C_{1}$, respectively, and $H_{x_{2}}=\left[I_{n_{3}}\otimes H_{4},H_{3}^{T}\otimes I_{n_{4}}\right]$ with dimension of $m_{3}\times n_{B}$ and $H_{z_{2}}=\left[H_{3}\otimes I_{n_{4}},\ I_{n_{3}}\otimes H_{4}^{T}\right]$ with dimension of $m_{4}\times n_{B}$ are the $X$-type and $Z$-type parity-check matrixes of toric code $C_{2}$, respectively. Let recall that $n_{A}=2n_{1}n_{2}=m_{1}+m_{2}$ and $n_{B}=2n_{3}n_{4}=m_{3}+m_{4}$ is the code length of $C_{1}$ and $C_{2}$, respectively.
According to Theorem 1 and $\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)=0$, the dimension $k$ of the 4D Chamon code is $k_{SV}+k_{TU}$, where $k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ and $k_{TU}=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$.
@@
Observing that the dimension $k_{SV}$ of $\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ is $\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\right)\times\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$, thus $k_{SV}=4\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$. Similarly, the dimension $k_{TU}$ of $\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$ is also $4\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$. Thus, the dimension $k$ of the 4D Chamon code is $k=k_{SV}+k_{TU}=8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, which completes the proof. ∎
For 4D toric code which is 4D homological product of two 2D toric codes, no matter what values $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ take, its dimension is always 6. While the dimension of 4D Chamon code, $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, is higher than that of 4D toric code, which means using the same two CSS codes, 4D XYZ product may construct codes with higher dimension than 4D homological product.
III.3 Logical operators and code distance of 4D XYZ product code
In this section, we study the logical operators and the code distance of 4D XYZ product code. First, we give out the general form of logical operators of the 4D XYZ product code in Theorem 2, then we compute their minimum weight in corollary 4.
@@ III.3 Logical operators and code distance of 4D XYZ product code
which will be used in the following proof.
Theorem 2 (The logical operators of 4D XYZ product code).
Giving two CSS codes $C_{1}$ and $C_{2}$, whose $X$-type and $Z$-type parity-check matrixes are $H_{x_{1}}$, $H_{z_{1}}$ and $H_{x_{2}}$, $H_{z_{2}}$, respectively, the logical operators of the corresponding 4D XYZ product code are divided into two types. The first type of logical operators are
@@
$\begin{bmatrix}X_{L_{2}}\\ X_{L_{3}}\\ Z_{L_{2}}\\ Z_{L_{3}}\end{bmatrix}=\begin{bmatrix}X(\textbf{a}_{1})\ &I\ &I\ &Y(\textbf{b}_{1})\ &I\\ I\ &Y(\textbf{a}_{2})\ &I\ &I\ &X(\textbf{b}_{2})\\ Y(\textbf{c}_{1})\ &Z(\textbf{d}_{1})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z(\textbf{c}_{2})\ &Y(\textbf{d}_{2})\end{bmatrix}$ (30)

### v3

@@ 3.2 Code dimension of 4D XYZ product code
Next, we first consider 4D XYZ product of two concatenated codes obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ ($n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd), and compute its code dimension. Consider the repetition code $C_{1}$ with length $n_{1}$ be the outer code and the code $C_{2}$ with length $n_{2}$ be the inner code, whose parity-check matrices are $H_{1}$ with dimension $\left(n_{1}-1\right)\times n_{1}$ and $H_{2}$ with dimension $\left(n_{2}-1\right)\times n_{2}$ respectively, the corresponding $X$-type and $Z$-type parity-check matrices $H_{x}$ and $H_{z}$ of the concatenated code obtained from $C_{1}$ and $C_{2}$ are as follows:
@@
Corollary 2 (The code dimension of 4D XYZ product concatenated code).
The code dimension $k$ of the 4D XYZ product of two concatenated codes obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ (where $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd) is 1.
@@
According to Theorem 1 and $\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)=1$, the code dimension $k$ of 4D XYZ product concatenated code is $1+k_{SV}+k_{TU}$, where $k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ and $k_{TU}=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$.
@@
For a concatenated code $C$ obtained from a pair of repetition codes, whose $X$-type and $Z$-type parity-check matrices are $H_{x}$ and $H_{z}$ as shown in Eq. (24) and Eq. (25) respectively, it is easy to prove that $\dim\left(\ker\left(\begin{bmatrix}H_{x}\\ H_{z}\end{bmatrix}\right)\right)=1$ and $\dim\left(\ker\left(\left[H_{z}^{T},H_{x}^{T}\right]\right)\right)=n-1-dim\left(row\left(\begin{bmatrix}H_{x}\\ H_{z}\end{bmatrix}\right)\right)=n-1-(n-1)=0$. The reason is that the undetectable $Y$-type error of code $C$ only contains one $Y$-type logical operator whose weight is code length. Thus, the code dimension $k$ of 4D XYZ product concatenated code is 1. ∎
Next, we consider the second instance of 4D XYZ product, 4D Chamon code, and compute its code dimension. Before computing it, we first prove Lemma 1 which will be used later.
Lemma 1 (The number of independent undetectable $Y$-type errors of two-dimensional (2D) toric code).
@@
Since 2D toric code encodes 2 logical qubits, there are 2 independent $Y$-type logical operators at most. When the toric code lattice is a square, these two independent $Y$-type logical operators are full diagonal of Pauli $Y$ operators from the upper right to the lower left and from the upper left to the lower right, respectively. In Fig. 4, we give an example of $Y$-type logical operators of 2D toric code with lattice size $3\times 3$.
@@
Any $Y$-type stabilizer must be generated by the product of some $X$-type and $Z$-type stabilizers both of which act on the same set of qubits. According to the topology of toric code, this can only be realized by selecting plaquettes ($Z$-type stabilizers) and vertices ($X$-type stabilizers) along the diagonal of $j\times j$ square lattices periodically. Fig. 5 gives an example of $Y$-type stabilizer of toric code with size of $j=4,\ k=6$.
Figure 5: A $Y$-type stabilizer of toric code with lattice size $j=4,\ k=6$. The red solid lines represent the qubits which the $Y$-type stabilizer acts on. The blue-grey plaquettes denote the selected $Z$-type stabilizers, while the top-right and bottom-left vertices indicate the selected $X$-type stabilizers.
When $j$ and $k$ are coprime, namely, $\gcd\left(j,k\right)=1$, it is impossible to select some $X$-type and $Z$-type stabilizers to generate a $Y$-type stabilizer, thus the number of $Y$-type stabilizers is zero. While in the case of $\gcd\left(j,k\right)=g\neq 1$, using the selection method as shown in Fig. 5, one can find $2\left(g-1\right)$ independent $Y$-type stabilizers by shifting to left or right, where the factor 2 comes from the left-right symmetry of toric code.
@@
Corollary 3 (The code dimension of the 4D Chamon code).
The code dimension $k$ of the 4D Chamon code, which is 4D XYZ product of two hypergraph product codes that are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$, is
@@
The hypergraph product of a pair of repetition codes is a 2D toric code[6]. Thus, given two toric codes which are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$, whose parity-check matrices are $H_{1}$, $H_{2}$, $H_{3}$ and $H_{4}$, respectively, by Definition 1, the corresponding stabilizer matrix S of the 4D Chamon code is as shown in Eq. (16), where $H_{x_{1}}=\left[I_{n_{1}}\otimes H_{2},H_{1}^{T}\otimes I_{n_{2}}\right]$ with dimension of $m_{1}\times n_{A}$ and $H_{z_{1}}=\left[H_{1}\otimes I_{n_{2}},\ I_{n_{1}}\otimes H_{2}^{T}\right]$ with dimension of $m_{2}\times n_{A}$ are the $X$-type and $Z$-type parity-check matrices of toric code $C_{1}$, respectively, and $H_{x_{2}}=\left[I_{n_{3}}\otimes H_{4},H_{3}^{T}\otimes I_{n_{4}}\right]$ with dimension of $m_{3}\times n_{B}$ and $H_{z_{2}}=\left[H_{3}\otimes I_{n_{4}},\ I_{n_{3}}\otimes H_{4}^{T}\right]$ with dimension of $m_{4}\times n_{B}$ are the $X$-type and $Z$-type parity-check matrices of toric code $C_{2}$, respectively. Recall that $n_{A}=2n_{1}n_{2}=m_{1}+m_{2}$ and $n_{B}=2n_{3}n_{4}=m_{3}+m_{4}$ is the code length of $C_{1}$ and $C_{2}$, respectively.
According to Theorem 1 and $\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)=0$, the code dimension $k$ of the 4D Chamon code is $k_{SV}+k_{TU}$, where $k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ and $k_{TU}=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$.
@@
Observing that the dimension $k_{SV}$ of $\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ is $\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\right)\times\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$, thus $k_{SV}=4\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$. Similarly, the dimension $k_{TU}$ of $\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$ is also $4\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$. Thus, the code dimension $k$ of the 4D Chamon code is $k=k_{SV}+k_{TU}=8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, which completes the proof. ∎
For 4D toric code which is 4D homological product of two 2D toric codes, its code dimension remains 6 regardless of the values of $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$. While the code dimension of 4D Chamon code, $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, is higher than that of 4D toric code, which means using the same two CSS codes, 4D XYZ product may construct codes with higher code dimensions than 4D homological product.
3.3 Logical operators and code distance of 4D XYZ product code
In this section, we study the logical operators and the code distance of 4D XYZ product code. First, we present the general forms of the logical operators in Theorem 2, followed by the proof of the upper bound of their minimum weight in corollary 4.
@@ 3.3 Logical operators and code distance of 4D XYZ product code
which will be used in the proof of Theorem 2.
Theorem 2 (The general forms of the logical operators of 4D XYZ product code).
Given two CSS codes $C_{1}$ and $C_{2}$, whose $X$-type and $Z$-type parity-check matrices are $H_{x_{1}}$, $H_{z_{1}}$ and $H_{x_{2}}$, $H_{z_{2}}$, respectively, the logical operators of the corresponding 4D XYZ product code are divided into two types. The first type of logical operators are
@@
$\begin{bmatrix}X_{L_{2}}\\ X_{L_{3}}\\ Z_{L_{2}}\\ Z_{L_{3}}\end{bmatrix}=\begin{bmatrix}X(\textbf{a}_{1}^{T})\ &I\ &I\ &Y(\textbf{b}_{1}^{T})\ &I\\ I\ &Y(\textbf{a}_{2}^{T})\ &I\ &I\ &X(\textbf{b}_{2}^{T})\\ Y(\textbf{c}_{1}^{T})\ &Z(\textbf{d}_{1}^{T})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z(\textbf{c}_{2}^{T})\ &Y(\textbf{d}_{2}^{T})\end{bmatrix}$ (31)